Can you please help me find:
center, vertices, covertices and foci of the ellipse x^2/4+y^2/9=16
Thank you so much!!

Answer by stanbon(48581)

center, vertices, covertices and foci of the ellipse x^2/4+y^2/9=16
------------------
x^2/64 + y^2/144 = 1
---------------------
a = 8 ; b = 12;
c = sqrt(12^2-8^2) = sqrt(80) = 4sqrt(5)
----------------------
Center: (0,0)
--------------------
vertices:
(0,12) ; (0,-12)
-----------
Foci:
(0,4sqrt(5))
(0,-4sqrt(5))
==============
Cheers,
Stan H.